OpenStudy (anonymous):

I am stuck for 2 days on this... Fourier transform of f(t) = sin(2t) / e^|t|

OpenStudy (anonymous):

f(t) = sin(2t) / e^t * U(t) + sin(2t) / e^(-t) * U(-t) = sin(2t) * e^-t * U(t) + sin(2t) * e^t * U(-t)

OpenStudy (anonymous):

Here's the definition I have \[\large F(\xi)=\mathcal{F}\left\{f(t)\right\}(\xi)=\int_{-\infty}^\infty f(t)e^{-i\xi t}~dt\] Yours may be different, so make the necessary corrections. You're given \(f(t)=\dfrac{\sin(2t)}{e^{|t|}}\). This may make the integration easier: \[\begin{align*} \sin(2t)&=2\sin t\cos t\\ &=2\left(\frac{e^{it}-e^{-it}}{2}\right)\left(\frac{e^{it}+e^{-it}}{2}\right)\\ &=\frac{e^{2it}-e^{-2it}}{2} \end{align*}\] So, the Fourier transform of the given function is \[\large{ \begin{align*}F(\xi)&=\frac{1}{2}\int_{-\infty}^\infty \frac{e^{2it}-e^{-2it}}{e^{|t|}}~dt\\ &=\frac{1}{2}\left[\int_{-\infty}^0 \frac{e^{2it}-e^{-2it}}{e^{-t}}~dt +\int_0^\infty \frac{e^{2it}-e^{-2it}}{e^t}~dt \right]\\ &=\frac{1}{2}\left[\int_{-\infty}^0 \left(e^{(1+2i)t}-e^{(1-2i)t}\right)~dt +\int_0^\infty \left(e^{-(1-2i)t}-e^{-(1+2i)t}\right)~dt \right] \end{align*}} \]

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

Hmm, I got total fourier transform = 0

OpenStudy (anonymous):

Oh wait, I left out the \(\xi\) term... Just a sec

OpenStudy (anonymous):

\[\large{ \begin{align*}F(\xi)&=\frac{1}{2}\int_{-\infty}^\infty \frac{e^{2it}-e^{-2it}}{e^{|t|}}~e^{i\xi t}~dt\\ \end{align*}}\]

OpenStudy (anonymous):

So the integral should be \[\large \frac{1}{2} \Bigg[ \int_{-\infty}^0 \left(e^{(1+2i)t}-e^{(1-2i)t}\right)e^{i\xi t}~dt\\ ~~~~~~~~~~~~~~~~~~~~\large+\int_0^\infty \left(e^{-(1-2i)t}-e^{-(1+2i)t}\right)e^{i\xi t}~dt \Bigg]\]

OpenStudy (anonymous):

I'll work out the first integral: \[\large\int_{-\infty}^0 \left(e^{(1+2i)t}-e^{(1-2i)t}\right)e^{i\xi t}~dt\\ \large\int_{-\infty}^0 \left(e^{(1+(2+\xi)i)t}-e^{(1-(2-\xi)i)t}\right)~dt\\ \large\left[ \frac{e^{(1+(2+\xi)i)t}}{1+(2+\xi )i}-\frac{e^{(1-(2-\xi)i)t}}{1-(2-\xi )i } \right]_{-\infty}^0\\ \large \left[ \frac{1}{1+(2+\xi )i}-\frac{1}{1-(2-\xi )i } \right] - \left[ 0-0 \right]\\ \large \frac{-2(2-\xi )i}{1+(2+\xi )^2} \] I hope I'm doing this right, it's been a while...

OpenStudy (anonymous):

The second integral would be done similarly. Finding a common denominator among the resulting terms should get you something like what WA is giving me: http://www.wolframalpha.com/input/?i=FourierTransform%5BSin%5B2*t%5D%2FExp%5B%7Ct%7C%5D%2Ct%2Cw%5D

OpenStudy (anonymous):

F(sin 2t) = pi/j * [$(w-2) - $(w+2) ] F ( sin(2t) * e^-t * U(t) ) = pi/j * [$(w-1) - $(w+3) ] similarily, F ( sin(2t) * e^t * U(-t) ) = -pi/j * [$(w-1) - $(w+3) ]

OpenStudy (anonymous):

we sure have different ways about this.. :)